OpenStudy (anonymous):

If a+b+c=1 , find maximum value of (ab+bc+ac). a,b,c are positive numbers.

OpenStudy (anonymous):

what is find maxm

OpenStudy (anonymous):

in my point of view. I would say a=0.33 b=0.33 and c=0.34

OpenStudy (anonymous):

then why a=0.1,b=0.4,c=0.5 is incorrect?

OpenStudy (anonymous):

Now i shall speak from my non-existent knowledge. Whatever number makes 1 is common sense. so it could be a=0.9 b=0.05 c=0.05

OpenStudy (anonymous):

but is you said MAX value. I would re-correct myself and say a= 0.3 b= 0.3 c=0.4

OpenStudy (anonymous):

see ques. carefully , i want maximum value of ab+bc+ac.

OpenStudy (anonymous):

try squaring both sides, there you can find ab+bc+ac

OpenStudy (anonymous):

A Mathematica solution is attached.

OpenStudy (anonymous):

@Luis_Rivera, you are wrong

OpenStudy (anonymous):

the maximum value you can achieve here only when a, b and c are equal. That's 3a=1, so each equals 1/3. Therefore, by substiting to equation that's given you will get 1/3. So, 1/3 is the maximum value

OpenStudy (anonymous):

using Lagrangian setup, you get the three conditions as: \[ b+c=\lambda\\ a+c=\lambda\\ b+a=\lambda\\ a+b+c=1 \] solving these four equations, we see that \[ b-a=0\implies a=b\\ c-a=0\implies a=c\\ a=b=c={\lambda\over2}\\ {3\lambda\over2}=1\implies\lambda={2\over3}\\ a=b=c={1\over3} \]